1. (25) An astronaut will fly a rocket in a cometary orbit given by f(x) = **-***Please find f and X-2 f in FACTORED FORM, So it 2 questions 1. And 2.
Answer #1

2-3 nonvert Asymptolog X-2 Minimum (313) flol= (4-5)/27-5)-(593; y=x-1 (x-2/2 - maximum (1,-1) vertical Asymptotes. y-interc